suppose an economy is described by the following equation:

Y = C + I + G + X - M

C = 14 + 0.60Yd

I = 20

G = 10

X = 15

M = 5 +0.1Y

T =20 +0.4 Y

(a) Find out the equilibrium value of income

(b)Calculate net export

(c) what is the value of export multiplier?

(d) if the equilibrium national income at full employment is N$202, what should be the increase in government spending or in exports to attain the full employment level of income?




1
Expert's answer
2022-03-11T08:32:15-0500

"Y = C + I + G + X - M"

"C = 14 + 0.60Yd"

"I = 20"

"G = 10"

"X = 15"

"M = 5 +0.1Y"

"T =20 +0.4 Y"


(a) Find out the equilibrium value of income

The expenditure approach method is given by;

"Y=C+I+G+(X-M)"


Therefore;

"Y=14+0.60(Y-(20 +0.4 Y)+20+10+(15-(20+0.4Y)"

"Y=0.60Y-12-0.24Y+30+15-20-0.4Y"

Put like terms together and simplify

"Y-0.60Y+0.24Y+0.4Y=8"

"1.04Y=8"

"Y=\frac{8}{1.04}"

"Y=7.69"


b)"net export=exports -imports"

"=15-(5+0.1Y)"

"=15-((5+(0.1\times7.69)"

"=15-0.769"

"=14.231"


c)"export multiplier=0.1Y"

"=0.1\times7.69"

"=0.769"


d)"202=0.60Y-12-0.24Y+30+15-20-0.4Y"

Put like terms together

"202+12-30-15+20=0.60Y-0.24Y-0.4Y"

"189=-0.04Y"

"Y=\frac{189}{-0.04}"

"Y=4725"

Y=$4,725